kdudley4:

Which expression is equivalent to 8a + 20? 4(2a + 20), 4(2a + 5), 4(4a + 16), 4(4a + 5)

Vocaloid:

use the distributive property A(B+C) = A*B+A*C try to apply this logic to each answer choice and see which one equals 8a + 20 when expanded let me know if you're having some difficulty or want to check your solution ^_^

kdudley4:

im having some difficulty

Vocaloid:

that's ok, let me try to explain in more detail ^_^ let's start with the first answer choice: 4(2a + 20) if we distribute the 4 we get 4 * 2a and 4 * 20 giving us 8a + 80 so that's not quite it can you try distributing 4(2a + 5) using the same line of logic?

kdudley4:

i see know, thank you (:

Vocaloid:

good so if we distribute 4(2a+5) we get 4*2a + 4*5 = 8a + 20 so 4(2a+5) is our solution

kdudley4:

can u help me with this problem too....... A theater company charges $7.50 per ticket for people of all age groups. If c represents the total number of children who bought tickets and a represents the total number of adults who bought tickets, which expression best represents the total money earned by the company from the sale of tickets?

Vocaloid:

well we can add the children and adults to get (c+a) since they're all being charged the same price the total revenue can be found by multiplying the ticket price by the total # of people in attendance so putting everything together something like 7.50(a+c)?

kdudley4:

thank you

kdudley4:

Zak and Ryan earn $60 per week for delivering pizzas. Zak worked for x weeks and earned an additional total bonus of $25. Ryan worked for y weeks. Which expression shows the total money, in dollars, that Zak and Ryan earned for delivering pizzas?

Vocaloid:

well, since both earn 60 per week delivering pizzas, we can multiply 60 * the number of weeks (in this case, x, and y) and then add on the 25 at the end any ideas what this might look like as an expression?

kdudley4:

60x + 25 + 60y

Vocaloid:

well done

kdudley4:

thank you

kdudley4:

Michelle did a survey in four butterfly parks and wrote the following observations: Butterfly Park Survey Park Number of Red Butterflies Total Number of Butterflies A 33 60 B 36 78 C 40 89 D 24 59 Based on the survey, which park had the greatest percentage of red butterflies?

Vocaloid:

percentage of red butterflies = 100*(red butterflies)/(total butterflies) so simply go through each answer choice, perform this calculation, and see which one gives the greatest result

kdudley4:

im still confused my teacher taught me that u have to multiply or divide it....or something like that. Do u have to multiply or divide it?

Vocaloid:

/ means divide

kdudley4:

ok thank u:p

kdudley4:

so the answer is park D

Vocaloid:

100*(24/59) = 40.7% but there's a park with a bigger percentage try starting with park A again and re-doing the calculations

kdudley4:

ok

kdudley4:

oh park C

Vocaloid:

100*(40/89) = 44.9% but this isn't the highest percentage hint: try park A. is it greater or less than 44.9%?

kdudley4:

oooohhhhhhhh, OK

kdudley4:

thank you

Vocaloid:

good, so park A has 55% which is greater than the others, so park A is the solution

kdudley4:

Nadia surveyed 80 students at her school and found that 0.55 of the students she surveyed said their favorite class is math. Another 15% of the students she surveyed reported that their favorite class is science. How many more students in the survey prefer math over science?